OpenStudy (anonymous):

find dy/dx and then put x = 1/4 y=x+1/sqrt{x}

OpenStudy (anonymous):

is anyoneee hereee pleasee replyy

OpenStudy (mathmale):

Is this the question you wanted to discuss?

OpenStudy (mathmale):

Assuming that you want to find the derivative of y=x+1/sqrt{x}: I'd strongly suggest that you re-write the function first, as\[f(x) = x +\frac{ 1 }{ \sqrt{x} }=x+x ^{\frac{ -1 }{ 2 }}\]

OpenStudy (mathmale):

Hope this helps you. Differentiate with respect to x, using the power rule. Simplify the results. Finally, evaluate this derivative at x = 1/4. f '(1/4) = ????